Vocaloid:

A person weighs 1050N when weighed in air and 50N when completely submerged in water. What is their body fat percentage? (given sp. gr of fat = 0.94 and sp. gr of other tissues = 1.1)

Vocaloid:

Please do not comment unless you are making a serious attempt at the problem.

imqwerty:

weight of vol of water displaced = buoyant force vol of water displaced = vol of body

imqwerty:

\(body~fat~fraction ~=~x\) \(other~tissue~fraction ~=1-x\) \(body~vol = V\) fraction of volume of fat = V*x fraction of volume of other tissue = V*(1-x) body weight = V*x*\(d_{fat}\) + V*(1-x)*\(d_{other ~tissue}\)

Vocaloid:

Thank you c;